OpenStudy (melissa_412):

Can someone help me with this quick 8th grade math question? I know it's late sorry but I am super sleepy and I need to get this done by tomorrow. Examine this system of equations. Which number can be multiplied by each equation so that when the two equations are added together, the y term is eliminated? 3/4x+2/3y=6 5/8x+5/6y=12 A. 15 times the first equation and 12 times the second equation B. 15 times the first equation and –12 times the second equation C. 30 times the first equation and –6 times the second equation D. 30 times the first equation and 6 times the second equation

OpenStudy (melissa_412):

@ganeshie8 @Opcode @@MightyChondrion

OpenStudy (melissa_412):

@MightyChondrion

OpenStudy (melissa_412):

Can you please help.?

OpenStudy (melissa_412):

@ParthKohli

OpenStudy (mightychondrion):

It's B. because if you try every solution they give it works out as follows: 15*(2/3y)=(30/3y)=10y and -12*(5/6y)=(-60/6y)=-10y and those two add together to cancel out.